In week 1 of the chemical equilibrium experiment, you and your lab team will create a calibration curve for FeSCN2+ which will relate the FeSCN2+ concentration (in M) to absorbance of light at a particular wavelength. Absorbance is plotted on the y-axis, and the FeSCN2+ concentration (in M) is plotted on the x-axis. If the equation for the regression line is y = 4193 x + 7.3×10−2, calculate the concentration of FeSCN2+ in a solution that has an absorbance of 0.291.
Answer #1

we have

y = 4193 x + 7.3×10^−2

slope = 4193

and

intercept = 7.3 * 10^-2

y axis ..............> absorbance

and

x axis .................> concentration.

when absorbance = 0.291 then

0.291 = 4193 x + 7.3 * 10^-2

or

x = 5.20 * 10^-5 M

thus

concentration of FeSCN2+ = 5.20 * 10^-5 M